a baker in 70 minutes iced 40 cupcakes and in 49 minutes iced 28 cupcakes. what is the constant of proposinalty



1
Expert's answer
2021-09-17T03:54:02-0400

Suppose "y=minutes" and "x=cupcakes"

So "y=kx+b"

Put "(40,70)" and "(28, 49)" into equation "y=kx+b"

So, "70=40k+b"

"49=28k+b"

Dividing equation: "70=40k+b" by 10 we get "7=4k+\frac{1}{10}b"

so "7-4k=\frac{1}{10}b"

Dividing equation: "49=28+b" by 7 we get "7=4k+\frac{1}{7}b"

so "7-4k=\frac{1}{7}b"

So "\frac{1}{10}b=\frac{1}{7}b"

So "b=0"

Substitute b into equation: "70=40k+b"

"70=40k"

"k=\frac{70}{40}=\frac{7}{4}"

"k=1.75"

Thus, the constant of proportionality is 1.75.